OpenStudy (ujjwal):

The solution of the equation \(\cos ^2 \theta +\sin \theta +1=0\), lies in the interval a)(-\(\pi\)/4, \(\pi\)/4) b)(\(\pi\)/4, 3\(\pi\)/4) c)(3\(\pi\)/4, 5\(\pi\)/4) d)(5\(\pi\)/4, 7\(\pi\)/4)

OpenStudy (anonymous):

cos^2 + sin^2 = 1

OpenStudy (ujjwal):

oh, I wrote the question wrong!

OpenStudy (ujjwal):

I have corrected it now!

OpenStudy (ujjwal):

@estudier

OpenStudy (anonymous):

Use cos^2 + sin^2 = 1 to rewrite, solve quadratic..

OpenStudy (anonymous):

change the \(\large cos^2\theta \rightarrow 1-sin^2\theta \). now you have a trignometric equation in quadratic form.. solve the quadratic.

OpenStudy (ujjwal):

I am being unable to solve the quadratic equation..

OpenStudy (anonymous):

You cannot solve x^2-x-2 where x= sin theta?

OpenStudy (ujjwal):

yeah, i can.. I was making a silly mistake...

OpenStudy (anonymous):

Happens to us all.....